OpenStudy (anonymous):

At a height of h meters, you can see V kilometers to the horizon. These numbers are related by the equation V=3.5sqrt h. A person can see 378 km to the horizon from an airplane window. How high is the airplane?

OpenStudy (unklerhaukus):

V=3.5√h h=(V/3.5)^2 =(378/3.5)^2 =11664m =11.7km